tax year 2020. Donald Chin works for Northwest Supplies. His rate of pay is $14.00 per hour, and he is paid 1½ times the regular rate for all hours worked in excess of 40 per week. During the last week of January of the current year, he worked 48 hours. Chin is married and claims three withholding allowances on his W-4 form. His weekly wages are subject to the following deductions:

  1. Employee federal income tax
  2. Social Security tax at 6.2%
  3. Medicare tax at 1.45%
  4. Health insurance premium, $85
  5. Credit union, $125
  6. United Way contribution, $10

Required

  1. What would be Chin’s (a) regular pay, (b) overtime pay, (c) gross pay, (d) and net pa?
Answer #1

Answer :-

Explanation :-

a) Regular pay = 40 × 14 = 560

b) overtime pay = 8×1 ×21 = 168

c) gross pay = 560 + 168 = 728

d)Net pay = gross pay - deductions

Deductions = employee federal tax (27) + social security tax (728 × 6.2% = 45.14) + medicare tax (728×1.45% = 10.556) + 85 + 125 + 10

= 302.696

Net pay = 728 - 302.696 = 425.304

Note:-

I assume the federal rate amount is 27 is the answer is slightly changes it varies.